### v1

# A unified model for long-period radio transients and white dwarf binary pulsars

Abstract
Long-period radio transients (LPTs) represent a recently uncovered class of Galactic radio sources exhibiting minute-to-hour periodicities and highly polarised pulses of second-to-minute duration. Their phenomenology does not fit exactly in any other class, although it might resemble that of radio magnetars or of white dwarf (WD) radio emitting binary systems. Notably, two LPTs with confirmed multi-wavelength counterparts have been identified as synchronised white dwarf–M dwarf binaries (polars). Meanwhile, systems such as AR Scorpii and J1912--44 exhibit short-period pulsations in a hr-tight orbit, with polarized radio emission proposed to be generated by the interaction of the WD magnetosphere with the low-mass companion wind.
Here we demonstrate that both LPTs and WD binary pulsars can be explained within a single geometric model in which radio emission is triggered when the magnetic pole of a rotating white dwarf intersects its companion’s wind in the binary orbital plane. We use a 36-year timing baseline to infer the orbital period and binary geometry solely from radio data of GPM J1839--10, the longest-active LPT known. The model naturally predicts its intermittent emission and double-pulse structure. Crucially, we show that the beat period between the spin and the orbit matches the observed pulse substructure and polarisation signatures, providing strong support for the model. Applying this model to the WD binary pulsar J1912--44  it can also reproduce the system emission and geometry. Our results place GPM J1839--10, and other LPTs in general and radio emitting WD binaries, at different stages of a continuum between intermediate and synchronised polars, suggesting a unified population of magnetic WD binaries driving coherent radio emission.
{affiliations}
International Centre for Radio Astronomy Research, Curtin University, Kent St, Bentley WA 6102, Australia
Institute of Space Sciences (ICE, CSIC), Campus UAB, Carrer de Can Magrans s/n, 08193, Barcelona, Spain
Institut d’Estudis Espacials de Catalunya (IEEC), Esteve Terradas 1, RDIT Building, 08860, Castelldefels, Spain
National Radio Astronomy Observatory, Socorro, NM 87801, USA
Australia Telescope National Facility, CSIRO, Space & Astronomy, Epping, New South Wales, Australia
White dwarf binary systems are increasingly recognised as hosts of exotic pulsar-like emission, challenging long-standing assumptions that coherent radio pulses are exclusive to neutron star systems[1, 2]. The discovery of long-period transients (LPTs), characterised by long periodicities (minutes to hours) , and strong polarisation, further deepens the mystery of how and where coherent radio emission originates [3, 4]. These intriguing systems are characterized by periodic radio pulses with high polarization (both linear and circular) and a high surface brightness pointing to a coherent radio emission process. However, the luminosities of the bright radio pulses in the $\sim$10 systems discovered thus far, exceed their rotational power, leaving uncertain the energy source. A strong magnetic field has been suggested to overcome this energy problem, suggesting at first the possibility of them being radio magnetars having perhaps had large fall-back at birth slowing down their periods [5, 6].
Two LPTs, ILT J1101$+$5521 [7] and GLEAM--X J0704--37 [8, 9], have recently been associated with a WD–M dwarf binary system via the detection of the orbital motion in the optical spectrum of the M companion star, at a period coincident with the radio pulse period. These systems might be hosting a WD with a spin synchronized with the orbital motion, hence belonging to the class of polars, despite not being in an accretion phase. In contrast, the AR Scorpii (Ar Sco) and J1912–44 radio emitting binary WDs are not synchronised with their companions; coherent radio pulses are observed at beat periods resulting from the interaction between the white dwarf spin and the orbital motion. The radio emission in these systems has been intriguing all along, and scenarios involving the dipolar losses of the white dwarf rotating field and the interaction with the M–dwarf companion wind have been proposed [10, 11, 12]
The phenomenological similarities between these sources invite a unified physical framework. We propose a geometric emission model in which radiation is generated when a WD’s magnetic pole sweeps across its companion star, injecting relativistic particles from the M dwarf’s wind into the WD’s magnetosphere [13, 14]. These particles are accelerated along magnetic field lines, producing a collimated radio beam aligned with the magnetic moment of the WD. The model incorporates orbital inclination, magnetic obliquity, and beam geometry, and predicts observable intensity variations as a function of the spin and orbital phase. Further intensity variations observed in the long-term radio outbursts of LPTs might be ascribed to the intrinsic variability of the M-dwarf wind.
Our model assumes a circular orbit with inclination angle $i$ between the orbital axis and the WD spin axis. The magnetic moment is tilted by angle $\alpha$ from the spin axis, and maximum radio emission is generated in the direction of the magnetic moment. Emission is modulated in two ways: (1) by the angle $\beta_{\text{MD}}$ between the magnetic pole and the M dwarf, governing the availability of particles via an assumed Gaussian function; and (2) by the angle $\beta$ between the magnetic pole and the Earth, determining visibility to observers. This model, although simplified, provides predictive power: emission is expected when both $\beta_{\text{MD}}$ and $\beta$ are small. The resulting phase-space of emission naturally leads to double-peaked pulse groups and intermittent visibility consistent with observations.
GPM J1839--10 is the longest lived LPT known, exhibiting a 22-minute period with sporadic pulses observed across 36 years of archival radio data [4]. In a coordinated campaign using MeerKAT, ASKAP, and VLA in 2024, we identified a higher-order periodicity: pulses cluster in double groups occurring every $\sim$8.75 hr. We interpret this as the orbital period $T=31482.4\pm 0.2$ s, which we are able to measure to such low uncertainty by taking advantage of the 36 year baseline.
Folding the data over this 8.75 hr modulation period and the previously reported 22-minute period ($P_{A}$)[4], as well as the derived beat period ($P_{B}$), reveals in Figure 1 that the pulse substructure aligns between pulses with $P_{B}=1265.2197\pm 0.0002$ s, rather than $P_{A}=1318.1957\pm 0.0002$ s. The substructure timescale of $\sim$50 s is consistent throughout the archival data, first discovered by autocorrelation between pulses [4], and folding on $P_{B}$ reveals further correlation and vertical alignment of the brightest peaks. This suggests that $P_{B}$ corresponds to the true spin period of the white dwarf (while $P_{A}$ is a spin-orbit beat period) inviting its interpretation as the cross-section of the radio beam.
The phase difference between the left and right pulse groups cannot be attributed to the difference in light travel time between orbital phases because that would be on the order of seconds for an orbit of this size, not minutes. Instead, we interpret the groups as arising from the intersections of the M dwarf’s highly inclined orbit with the cone traced out by the WD’s magnetic pole. This explains both the intermittency and the phase delay between pulse groups. Our geometric model reproduces the observed dynamic pulse profile as shown in Figure 2 when fitted using a Markov chain Monte Carlo method, with best-fit parameters indicating a near-perpendicular orbital inclination $i=100.1\pm 0.6^{\circ}$. The modelled pulse groups are the regions where the $\beta_{\text{MD}}$ contours cut into the central column of the $\beta$ colour map — where the Earth, MD, and WD pole are sufficiently aligned. No good fit exists for $P_{A}$ as the spin period.
[…]
Figure 2: Modelled dynamic pulse profile of GPM J1839--10. At top left, the flux density predicted by the model $I_{\text{pred}}$ using the best-fit parameters found using MCMC is overlaid on the real pulse profiles. At top right, the colourmap is the LOS-beam angle $\beta$ as a function of spin-orbit phase, and the contours are the beam-MD angle $\beta_{\text{MD}}$. $I_{\text{pred}}$ is greatest where $\beta$ and $\beta_{\text{MD}}$ are the smallest. The bottom panel is a full orbit recorded by ASKAP and the associated model prediction. The vertical lines are spaced by the spin period.
The emission is consistently highly linearly polarised with occasional 90 jumps in polarisation angle (PA) to orthogonal polarisation modes (OPMs)[4]. However, Figure 1 reveals that the two pulse groups which appear at different orbital and spin phases are dominated by different OPMs, with a similar amount of variation in each group ($\sigma_{\text{PA}}\approx{\sim}10^{\circ}$). The PAs of individual pulses, however, may evolve more gradually and systematically across spin phase [15] than can be seen in the point clusters in the figure. It can further be seen that there is a small downturn in the PAs at around spin phase $-0.05$ (the trailing edge of the first group), as well as a similar upturn in the PAs at spin phase $0.05$ (the leading edge of the second group). Despite the possibility that pulses arriving between the two groups could potentially connect the PAs smoothly across intermediate spin/orbital phases, the flatness of the PAs of the two groups themselves, as well as the occasional $90^{\circ}$ jump within each group, lead us to prefer an OPM interpretation over a rotating vector model one[16].
Although not well understood, OPMs in pulsars are thought to be magnetospheric in origin, possibly due to birefringent properties of the co-rotating plasma[17]. We advance a similarly broad interpretation here: the OPMs are dominated by the average magnetic properties of the plasma through which the WD beam propagates along our line of sight. The consistency with which each pulse group exhibits the same dominant OPM support our geometric model, with the line of sight intersecting different physical sections of the intrabinary magnetic field during the two pulse groups. Both the single pulse PA variations and the occasional OPM jump even within a pulse group may be due to turbulence in the intrabinary wind.
Other spectral characteristics (dispersion measure, rotation measure, spectral shape) remain constant across the orbit, suggesting that variability arises primarily from geometric effects rather than intrinsic source changes on a short timescale, or propagation effects. On longer timescales (a few orbits) variability is primarily driven by MD wind activity.
To describe the system geometry, we place the WD at the origin with $\hat{z}$ as the WD spin axis, and the orbital axis inclined about $\hat{y}$. Fitting the model for GPM J1839--10, we find an orbital inclination $i=100.1\pm 0.6^{\circ}$ and magnetic obliquity $\alpha=52.1\pm 0.4^{\circ}$ with respect to $\hat{z}$. Our line of sight has azimuth $\phi_{0}=179.29\pm 0.04^{\circ}$ measured from $\hat{x}$ and zenith $\zeta=61\pm 2^{\circ}$ measured from $\hat{z}$. The radio beam has an opening angle of $W_{\text{spin}}=65\pm 3^{\circ}$ and becomes active when the MD is within $W_{\text{orb}}=70\pm 1^{\circ}$ of the WD magnetic moment $\hat{\mu}$. In Figure 3 we show the derived beam angle and orbit of the system together with critical radii to understand the emission state of the binary. The derived orbital period and WD spin period imply that: 1) the whole system is inside the WD light cylinder, 2) the companion star, if assumed to be an M-dwarf (as for the other two LPTs), is not filling its Roche-Lobe, and 3) depending on the assumed M-dwarf mass loss rate ($\sim 10^{-14}$ M yr-1 [18]), the WD Alfvén radius might be inside the corotation radius, implying that during stellar flares the system may undergo accretion episodes (possibly showing transient X-ray emission as observed in the LPT ASKAP J1832-0911 [19]), while being in propeller most of the time [20, 14]. Furthermore, the system geometry clearly shows how the rotating WD magnetosphere is naturally engulfed by the MD wind, possibly being the seed particles that are accelerated by the sweeping magnetic field lines generating the pulsed synchrotron radio emission [21, 22, 23].
Figure 3: Diagram of the binary system, in a moving reference frame centred on the WD. At left is a not-to-scale diagram of the geometric parameters. In green is WD with its magnetic moment vector. In red, the MD with its orbital path. The blue vector points towards Earth. At right is a to-scale projection of GPM J1839--10 on the y-z plane, with orbital phases marked. The cone traced by $\hat{\mu}$ crosses the y-z plane at the green full and dashed lines. The shaded red, orange, and cyan regions cover the ranges of $R_{\text{MD}}$, $R_{rl}$, and $R_{c}$ respectively, given the MD mass is $M_{\text{MD}}\in[0.14,0.5]M_{\odot}$ and the WD mass is $M_{\text{WD}}\in[0.6,1.2]M_{\odot}$. The green sector and solid red line are $W_{\text{spin}}$ and $W_{\text{orb}}$ respectively. The small and large black dashed circles are the Alfvén radii for the minimum and maximum WD masses respectively, assuming a stellar mass loss rate of $\dot{M}_{\text{MD}}$ = $10^{-14}$ M yr-1 [18].
To test the generality of the model, we applied it to 7.5 hours of MeerKAT observations of the WD binary pulsar J1912--44. Folding the data over the 5.3 minute spin and 4.03 hour orbital periods [2] forms the vertically aligned dynamic pulse profile in Figure 6. We used a prior of $i=59\pm 6^{\circ}$ based on the inclination angle constraints as measured in the optical band [14]. Given the prior, we find the following constraints on the geometry: $i\in[53,65]^{\circ}$, $\alpha\in[42,70]^{\circ}$, $\phi_{0}\in[-48,-17]^{\circ}$, $\zeta\in[39,73]^{\circ}$, $W_{\text{spin}}\in[20,60]^{\circ}$, and $W_{\text{orb}}\in[67,90]^{\circ}$. These results agree broadly with the earlier proposed geometry [14], but with only one pulse group the geometrical parameters are less constrained than for GPM J1839--10.
A critical component of the model is that the direction of the radio beam is locked to the WD rotating frame in the direction of the magnetic moment. If this is not the case it becomes difficult to explain the paired pulse group morphology, and it becomes unclear why the pulse substructure would be related to the spin phase. This is difficult to reconcile with models which place the emission site at or near the companion like the electron cyclotron maser emission (ECME) model [24] proposed for ILT J1101$+$5521 or the bow shock model [12] proposed for AR Sco. One scenario which allows for an emission region locked to the WD involves relativistic electrons accelerating from the companion towards the WD along the closed magnetic field lines when the companion star’s surface is heated by magnetic interaction [25, 23]. The electrons emit synchrotron radiation as they reach magnetic mirror points on their way to the WD pole [23]. This idea has been shown to reproduce the dynamic pulse profile of Ar Sco with spin and orbit modulated components similar to our model [10].
In any case, the emission mechanism must differ substantially from the common radio pulsar curvature radiation process, a testable prediction of which is the beam opening angle confined by the last open field lines at the pole. Assuming that the emission site is within the orbital radius ($r_{em}\leq A\leq 2.55R_{\odot}$) places a generous limit on the half-opening angle of $\rho\leq 15^{\circ}$, inconsistent with the best-fit half-opening angle $W_{\text{spin}}=65\pm 3^{\circ}$. Further separating GPM J1839--10 from neutron stars is the lack of PA sweep typically associated with the rotating vector model, although this is not a unifying feature of WD pulsars [26].
Future work will include applying the model to the growing number of LPTs, which will not only test it, but serve to constrain the distribution of their geometries. Of particular interest is the 6.45 h LPT ASKAP J183950.5--075635.0 [27] which has timing variation and pulse profile evolution phenomenologically similar to GPM J1839--10 but with an interpulse, extending the geometric parameter space. Also of interest would be J1755--2527, recently confirmed to be a 1.16 hour intermittent LPT [28], the emission-state-switching 54 minute LPT ASKAP J1935+2148 [29] and CHIME J0630$+$25 [30], whose 421 s period makes it an interesting counterpart to J1912--44.
Our results demonstrate that LPTs like GPM J1839--10 can be understood within the same emission geometry framework as radio emitting WD binary pulsars such as J1912--44 and AR Sco. The diversity in the observed periods, pulse morphologies, and polarisation states arises from variations in orbital synchronisation and viewing geometry. This implies that coherent radio emission from magnetic WD binaries is more common and varied than previously thought, potentially forming a new class of compact binary radio emitters. However, the possible evolutionary connection between these systems remains debated, and we cannot exclude them being at different evolutionary stages but with common properties and geometries allowing them to emit bright polarized radio emission. With enough samples, we expect their evolutionary stages to emerge, and the tools developed in this work will enable insight into binary WD and magnetic field evolution in a way difficult to select from optical surveys alone.
Acknowledgements
This research is supported by an Australian Government Research Training Program (RTP) Scholarship. N.H.-W. is the recipient of an Australian Research Council Future Fellowship (project number FT190100231). N.R. is supported by the European Research Council (ERC) via the Consolidator Grant “MAGNESIA” (No. 817661) and the Proof of Concept “DeepSpacePulse” (No. 101189496), by the Catalan grant SGR2021-01269 (PI: Graber/Rea), the Spanish grant ID2023-153099NA-I00 (PI: Coti Zelati), and by the program Unidad de Excelencia Maria de Maeztu CEX2020-001058-M.
This scientific work uses data obtained from Inyarrimanha Ilgari Bundara, the CSIRO Murchison Radio-astronomy Observatory. We acknowledge the Wajarri Yamaji People as the Traditional Owners and native title holders of the Observatory site. CSIRO’s ASKAP radio telescope is part of the Australia Telescope National Facility (https://ror.org/05qajvd42). Operation of ASKAP is funded by the Australian Government with support from the National Collaborative Research Infrastructure Strategy. ASKAP uses the resources of the Pawsey Supercomputing Research Centre. Establishment of ASKAP, Inyarrimanha Ilgari Bundara, the CSIRO Murchison Radio-astronomy Observatory and the Pawsey Supercomputing Research Centre are initiatives of the Australian Government, with support from the Government of Western Australia and the Science and Industry Endowment Fund.
The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc.
We acknowledge Ben Stappers, Ewan Barr, and Yunpeng Men for contributing to the MeerKAT proposal. We acknowledge Scott Hyman, Tracy Clarke, and Simona Giacintucci for contributing to the VLA proposal. We thank Ingrid Pelisoli and Simone Scaringi for useful discussions on white dwarf evolution.
Author contributions
Cs.H performed the orbital timing, spin/beat period analysis, polarisation analysis, developed the geometric model, and contributed to the discussion in the manuscript. N.R contributed to the formulation of the model and its physical interpretation, and contributed to the discussion in the manuscript. N.H.-W led the proposals for the radio data toward GPM J1839--10, imaged and produced dynamic spectra from the radio data used in this work, and contributed to the formulation of the model and discussion in the manuscript. S.M contributed to the formulation of the model, and contributed to the discussion in the manuscript. R.P generated the observing scripts and calibrated the data for the VLA observations. E.L processed the ASKAP data and contributed to the polarisation analysis.
Competing interests
The authors declare no competing interests.
Data availability
Data supporting the findings of this study are available from the corresponding author upon request.
Code availability
The code used for the geometric model and MCMC fitting is available at
https://github.com/CsanadHorvath/GeometricLPT.git.
An interactive web visualisation is available at
https://www.desmos.com/3d/6bisspujwe.