### v2

# A binary model of long period radio transients and white dwarf pulsars

Abstract
Long-period radio transients (LPTs) represent a recently uncovered class of Galactic radio sources exhibiting minute-to-hour periodicities and highly polarised pulses of second-to-minute duration. Their phenomenology does not fit exactly in any other class, although it might resemble that of radio magnetars or white dwarf (WD) radio emitting binary systems. Notably, two LPTs with confirmed multi-wavelength counterparts have been identified as WD – M dwarf binaries. Meanwhile, systems such as AR Scorpii and J1912-44 exhibit short-period pulsations in hrs-tight orbits, with polarised radio emission proposed to be generated by the interaction of the WD magnetosphere with the low-mass companion wind.
in which radio emission is triggered when the magnetic axis of a rotating WD intersects its companion’s wind in the binary orbital plane. We use a 36-year timing baseline to infer the orbital period and binary geometry from radio data alone. The model naturally predicts its intermittent emission and double-pulse structure. Crucially, we show that the beat period between the spin and the orbit matches the observed pulse substructure and polarisation signatures, providing strong support for the model. Applying it to the WD pulsar J1912-44, it successfully reproduces the emission profile and geometry as well. R1.1,R3.1]Our results suggest analogous emission-site geometries in these related classes of binary system — a possibility we extend to the broader LPT / WD pulsar population.
{affiliations}
International Centre for Radio Astronomy Research, Curtin University, Kent St, Bentley WA 6102, Australia
Institute of Space Sciences (ICE, CSIC), Campus UAB, Carrer de Can Magrans s/n, 08193, Barcelona, Spain
Institut d’Estudis Espacials de Catalunya (IEEC), Esteve Terradas 1, RDIT Building, 08860, Castelldefels, Spain
National Radio Astronomy Observatory, Socorro, NM 87801, USA
Australia Telescope National Facility, CSIRO, Space & Astronomy, Epping, New South Wales, Australia
1 Introduction
White dwarf binary systems are increasingly recognised as hosts of exotic pulsar-like emission, challenging long-standing assumptions that coherent radio pulses are exclusive to neutron star systems[1, 2]. The discovery of long-period transients (LPTs), characterised by their long periodicities (minutes to hours) compared to the bulk of the pulsar population, further deepens the mystery of how and where pulsed coherent radio emission originates [3, 4]. These intriguing systems are characterised by periodic radio pulses with high polarisation (both linear and circular) and a high surface brightness, pointing to a coherent radio emission process. However, the luminosities of the bright radio pulses in R1.4]many of the $\sim$12 systems discovered thus far exceed their spin-down power given typical pulsar-like assumptions, leaving uncertain the energy source. A strong magnetic field has been suggested to overcome this energy problem, suggesting at first the possibility of them being radio magnetars having perhaps had large fall-back at birth slowing down their periods [5, 6].
Two LPTs, ILT J1101++55 [7] and GLEAM-X J0704-37 [8, 9], have recently been associated with WD – M dwarf (MD) binary systems via the detection of the orbital motion in the optical spectrum of the M companion star, at a period coincident with the radio pulse period. These two systems might be hosting a WD with a spin synchronised with the orbital motion, despite not being in an accretion phase. R1.2,R1.5]We note, however, that the radio period need not necessarily equal the WD spin period. , which are not synchronised with their companions; coherent radio pulses are observed at the WD spin period and/or beat periods resulting from the interaction between the WD spin and the orbital motion, modulated on the orbital period with the brightest pulses arriving at a particular orbital phase. The radio emission in these systems has been intriguing all along, and scenarios involving the dipolar losses of the WD rotating field and the interaction with the MD companion wind have been proposed [12, 13, 14, 15, 16].
GPM J1839-10 is the longest-lived LPT known, exhibiting a 22-minute period with sporadic pulses observed across 36 years of archival radio data [4]. R1.8]The 30 to 300 s duration broadband radio pulses exhibit both linear and circular polarisation and quasiperiodic substructure. R1.4]Its original classification as a neutron star was based on its coherent radio luminosity ($\sim 3.5\times 10^{31}$ erg s-1 peak isotropic luminosity) but the period derivative $\dot{P}\lesssim 3.6\times 10^{-13}$ s s-1 [4] puts it in tension with the most generous pulsar emission models [18]. GPM J1839-10 also displays short duration ($\sim$20 ms) narrow-band sub-pulse features[19], suggesting multiple processes.
Motivated by the detection of a longer periodicity in observations of GPM J1839-10, we put forward that it is also a WD with a low-mass stellar companion, and propose a geometric model in which radio emission from the WD polar region is modulated by the relative alignment between the WD magnetic moment and the companion star’s position. We proceed on the premise that the aforementioned alignment determines the abundance of particles injected from the companion’s wind into the WD’s magnetosphere [20, 10]. These particles are accelerated along magnetic field lines, producing a collimated radio beam aligned with the magnetic moment of the WD. The model incorporates orbital inclination, magnetic obliquity, and beam geometry, and predicts observable intensity variations as a function of the spin and orbital phase. Further intensity variations observed in the long-term radio outbursts of LPTs might be ascribed to the intrinsic variability of the companion’s wind.
R1.1,R2.1,R3.1]Similar models have been proposed for J1912-44 [10] and AR Sco [12, 13], the latter of which reproduces AR Sco’s more complex pulse morphology with spin and orbit modulated components resembling GPM J1839-10. The phenomenological similarities between these sources and the geometric similarity invite an analogous physical framework.
2 Results
In a coordinated campaign using MeerKAT, ASKAP, and VLA in 2024, we obtained a near-contiguous 36 hour track of GPM J1839-10, in which we identified a higher-order periodicity: pulses cluster in paired groups occurring every $\sim$8.75 hr. We interpret this as an orbital period $P_{\text{orb}}=31482.4\pm 0.2$ s, which we are able to measure to such low uncertainty by taking advantage of the 36 year baseline (see subsection 4.3).
Folding the data over this 8.75 hr modulation period and the previously reported[4] 22-minute period ($P_{1}$), as well as the beat period ($P_{2}$) between $P_{1}$ and $P_{\text{orb}}$, creates the dynamic pulse profiles in Figure 1, revealing that the pulse substructure aligns between pulses with $P_{2}=1265.2197\pm 0.0002$ s, rather than $P_{1}=1318.1957\pm 0.0002$ s. The substructure timescale of $\sim$50 s is consistent throughout the archival data, first discovered by autocorrelation between pulses [4], and folding on $P_{2}$ reveals further correlation and vertical alignment of the brightest peaks. This suggests that $P_{2}$ corresponds to the true spin period of the white dwarf (while $P_{1}$ is a spin-orbit beat period) inviting its interpretation as the cross-section of the radio beam.
[…]
2.1 Modelling
The phase difference between the left and right pulse groups in Figure 1 cannot be attributed to the difference in light travel time between orbital phases because that would be on the order of seconds for an orbit of this size, not minutes. Instead, we consider a WD with a magnetic moment $\vec{\mu}$ misaligned from its spin axis $\hat{z}$ by angle $\alpha$ and a companion low-mass star (assumed to be an MD for simplicity) in a circular orbit whose axis is inclined by angle $i$ from $\hat{z}$ about $\hat{y}$ (see Figure 3). When the spin and orbital axes are misaligned, there are two orbital phases at which the rotating magnetic moment vector intersects the MD, which we suggest correspond to the observed pulse groups. Therefore, we model the observed radio emission as modulated by
  1. 1.
    a Gaussian function of the angle $\beta_{\text{MD}}$ between the WD magnetic moment and the MD, governing the availability of the MD’s wind at the WD pole, and
  2. 2.
    a Gaussian function of the angle $\beta$ between the WD magnetic moment and the line of sight (LOS), which is the cross-section of the radio beam parallel to the WD magnetic moment.
This model, although simplified, provides predictive power: emission is expected when both $\beta_{\text{MD}}$ and $\beta$ are small. The resulting phase-space of emission naturally leads to double-peaked pulse groups and intermittent visibility consistent with observations.
Fitting the model for GPM J1839-10 using a Markov chain Monte Carlo method, we find a near-perpendicular orbital inclination $i=100.1\pm 0.6^{\circ}$ and magnetic obliquity $\alpha=52.1\pm 0.4^{\circ}$. Our LOS has azimuth $\phi_{0}=179.29\pm 0.04^{\circ}$ and zenith $\zeta=61\pm 2^{\circ}$. The radio beam has an opening angle of $W_{\text{spin}}=65\pm 3^{\circ}$ and becomes active when the MD is within $W_{\text{orb}}=70\pm 1^{\circ}$ of $\hat{\mu}$. Figure 2 is the predicted dynamic pulse profile, alongside a map of $\beta$ and $\beta_{\text{MD}}$, and the lightcurve of a representative orbit. Intuitively, the modelled pulse groups are the regions where the $\beta_{\text{MD}}$ contours cut into the central column of the $\beta$ colour map — where the Earth, MD, and WD pole are sufficiently aligned. The model also predicts weak emission in the central spin phase $\sim$ orbit phase $\sim$ 0 region, where we indeed see sporadic dim pulses at about the substructure timescale. No good fit exists for $P_{1}$ as the spin period. In Figure 3 we show the derived beam angle and orbit of the system together with critical radii on a 2D projection.
To test the generality of the model, we applied it to previously published [2] 7.5 hours of MeerKAT observations of the WD binary pulsar J1912-44. Folding the data over the 5.3 minute spin and 4.03 hour orbital periods forms the vertically aligned dynamic pulse profile in Figure 4. We used a prior of $i=59\pm 6^{\circ}$ based on the inclination angle constraints as measured in the optical band [10]. Given the prior, we find the following constraints on the geometry: $i\in[53,65]^{\circ}$, $\alpha\in[42,70]^{\circ}$, $\phi_{0}\in[-48,-17]^{\circ}$, $\zeta\in[39,73]^{\circ}$, $W_{\text{spin}}\in[20,60]^{\circ}$, and $W_{\text{orb}}\in[67,90]^{\circ}$. These results agree broadly with the earlier proposed geometry [10], but with only one pulse group the geometrical parameters are significantly less constrained than for GPM J1839-10. The degeneracy of the J1912-44 parameters suggests that the vertical single-pulse-group morphology should be the most commonly observed among WD pulsars and LPTs.
Figure 2: Modelled dynamic pulse profile of GPM J1839-10. At top left, the flux density predicted by the model $I_{\text{pred}}$ using the best-fit parameters found using MCMC is overlaid on the real pulse profiles. At top right, the colourmap is the LOS-beam angle $\beta$ as a function of spin-orbit phase, and the contours are the beam-MD angle $\beta_{\text{MD}}$. The bottom panel is a full orbit recorded by ASKAP normalised to 1 GHz and the associated model prediction. The vertical lines are spaced by the spin period.
Figure 3: Diagram of the binary system, in a moving reference frame centred on the WD. At left is a not-to-scale diagram of the geometric parameters. In green is WD with its magnetic moment vector. In red, the MD with its orbital path. The blue vector points towards Earth. At right is a to-scale projection of GPM J1839-10 on the y-z plane, with orbital phases marked. The cone traced by $\hat{\mu}$ crosses the y-z plane at the green full and dashed lines. The shaded red, orange, and cyan regions cover the ranges of $R_{\text{MD}}$, $R_{rl}$, and $R_{c}$ respectively, given the MD mass is $M_{\text{MD}}\in[0.14,0.5]M_{\odot}$ and the WD mass is $M_{\text{WD}}\in[0.6,1.2]M_{\odot}$. The green sector and solid red line are $W_{\text{spin}}$ and $W_{\text{orb}}$ respectively. The small and large black dashed circles are the Alfvén radii for the minimum and maximum WD masses respectively, assuming a stellar mass loss rate of $\dot{M}_{\text{MD}}$ = $10^{-14}$ M yr-1 [21].
Figure 4: Modelled dynamic pulse profile for J1912-44.
3 Discussion
R2.2]Strengthening the case for a WD LPT progenitor beyond the thus-far mentioned are the LPTs ASKAP J1448-68 [22] and CHIME/ILT J1634++44 [23, 24], both of which have been shown to have optical spectral properties consistent with a WD. Furthermore, hints of accretion in the former and pulse arrival time properties in the latter imply possible binary systems.
For GPM J1839-10, Figure 3 shows how the rotating WD magnetosphere is naturally engulfed by the MD wind, possibly being the seed particles that are accelerated by the magnetic field lines generating the pulsed radio emission [25]. It further shows that: 1) the whole system is inside the WD light cylinder, 2) the companion star, if assumed to be an MD (as for the other two LPTs), is not filling its Roche-Lobe, and 3) depending on the assumed M-dwarf mass loss rate ($\sim 10^{-14}$ M yr-1 [21]), the WD Alfvén radius might be inside the corotation radius, implying that during stellar flares the system may undergo accretion episodes (possibly showing transient X-ray emission as observed in the LPT ASKAP J1832-0911 [26], R1.4,R3.6]although the X-ray luminosity limit of $L_{X}<2\times 10^{33}$ erg s-1[4] for GPM J1839-10 does not preclude some level of accretion[27]), while being in propeller most of the time [28, 10]. ASKAP J1448-68 has also been observed to emit X-rays[22], and its multi-wavelength properties are consistent with a magnetic white dwarf, possibly an accreting binary system. In general, it seems that significant accretion may not be a necessary component of the radio emission mechanism (whatever it may be) since we both do [26, 22, 23, 24, 2] and do not [7, 8, 1, 17] see signs of such across the LPT / WD pulsar population.
1.8,R2.1,3.1]It has been shown [12, 13] that for AR Sco’s radio emission, like for that of GPM J1839-10 in this work, the pulse morphology is best modelled with a beam direction locked to the WD rotating frame, as opposed to the binary frame, and modulated by the companion’s location in that frame. One may expect the geometric features between the discussed LPTs and WD pulsars to be similar simply because both originate from binary systems, but the comparison is stronger than that. These geometric constraints give us the important clue that the emission direction is dominated by the WD’s local field, while powered by interaction with the MD. If this is not the case it becomes difficult to explain the paired pulse group morphology, and it becomes unclear why the pulse substructure would be related to the spin phase. R1.9]This is difficult to reconcile with models like electron cyclotron maser emission (ECME) [29] or models which place the emission site near the MD [15, 30].
A scenario which has been suggested for AR Sco [1, 25] which allows for an emission region locked to the WD involves relativistic electrons accelerating from the companion towards the WD along the closed magnetic field lines when the companion star’s surface is heated by magnetic interaction. The electrons emit synchrotron radiation as they reach magnetic mirror points on their way to the WD pole. This idea was used successfully in the earlier mentioned geometric model [12] to predict polarimetric properties of AR Sco. R1.8]The broadband nature and spectral shape of GPM J1839-10 could be consistent with this, but the radio luminosity is orders of magnitude brighter. Despite the possibility that the radio emission from AR Sco and GPM J1839-10 are powered by different mechanisms, we highlight the consistent interpretation in the two systems that the emission is modulated as the source of particles (i.e. the MD) moves along its orbit. If the mechanisms differ then it is likely in the acceleration mechanism of said particles.
R1.9]Alternative forms of ECME in LPTs have been proposed and remain viable, such as relativistic ECME [14] or loss-cone-driven maser (LCDM) [16], which produce beamed emission aligned with the local magnetic field of the WD. R3.5]Less luminous ($L\lesssim 10^{25}$ erg s-1) radio emission from a related class of WD binaries, (magnetic) cataclysmic variables (CV / mCV), has been observed[31, 32]. If this emission is ECME, consistent with their typical circular polarisation [33] (not a requirement in the relativistic case), then there could be a natural spectrum from non-relativistic to relativistic regimes.
R1.6,R1.9,R3.2]In the case of ILT J1101++55, relativistic ECME was proposed to be powered by unipolar induction, in which an electric field is induced in the MD by the penetrating WD magnetic field, leading to the establishment of a current loop along the field lines connecting the two. However, this requires that the orbital separation is sufficiently small that the WD magnetic field dominates at the location of the MD. According to recent theoretical work [16], when the orbit is greater than $P_{\text{orb}}\sim 3$ hr, the MD magnetic field is strong enough to prevent the WD field from penetrating its surface, and we transition from the unipolar induction phase to the magnetospheric interaction phase. In this phase, radio emission occurs along magnetic field lines connecting the WD to a magnetic reconnection region between the two, rather than the MD itself. The orbit of GPM J1839-10 would place it in the magnetospheric interaction phase, and we calculate that the observed luminosity can be reached without requiring a strongly magnetic WD.
In any case, the emission mechanism must differ substantially from the common radio pulsar curvature radiation process, a testable prediction of which is the beam opening angle confined by the last open field lines at the pole. Assuming that the emission site is within the orbital radius ($r_{em}\leq A\leq 2.55R_{\odot}$) places a generous limit on the half-opening angle of $\rho\leq 15^{\circ}$, inconsistent with the best-fit half-opening angle $W_{\text{spin}}=65\pm 3^{\circ}$.
3.1 Polarisation
The emission of GPM J1839-10 is consistently highly linearly polarised (up to 100%) with occasional 90 jumps in polarisation angle (PA) to orthogonal polarisation modes (OPMs). However, Figure 1 reveals that the two pulse groups which appear at different orbital and spin phases are dominated by different OPMs, with a similar amount of variation in each group ($\sigma_{\text{PA}}\approx{\sim}10^{\circ}$). The PAs of individual pulses evolve more gradually and systematically across spin phase than can be seen in the point clusters in Figure 1 (see [19]), but we focus on the average behaviour. Both the single pulse PA variations and the occasional OPM jumps within a pulse group may be due to turbulence in the intrabinary wind, R1.10]but why one OPM should be preferred at a particular spin/orbit phase is unclear. Other spectral characteristics (dispersion measure (DM), rotation measure (RM), spectral shape) remain constant across the orbit, although DM and RM variations can be suppressed by the surrounding magnetospheric plasma.
It can further be seen that there is a small downturn in the PAs at around spin phase 0.05-0.05 (the trailing edge of the first group), as well as a similar upturn in the PAs at spin phase $0.05$ (the leading edge of the second group). Despite the possibility that dim pulses arriving between the two groups could potentially connect the PAs smoothly across intermediate spin–orbit phases (with greater sensitivity), the flatness of the PAs of the two groups themselves, as well as the occasional $90^{\circ}$ jump within each group, lead us to prefer an OPM interpretation over a rotating vector model[34].
Although not well understood, OPMs in pulsars are thought to be magnetospheric in origin, possibly due to birefringent properties of the co-rotating plasma[35]. We advance a similarly broad interpretation here: R1.10]the linear polarisation is dominated by the average magnetic properties of the plasma through which the WD beam propagates along the LOS, switching the visible polarisation mode with spin phase. If the mechanism behind the pulse group PA separation and the short duration OPMs are related, then further study of GPM J1839-10 may give insights into OPMs more broadly.
4 Conclusion
Future work will include applying the model to the growing number of LPTs, which will not only test it, but serve to constrain the distribution of their geometries. Of particular interest is the P=P=6.45-h LPT ASKAP J183950.5-075635.0 [36] which has timing variation and pulse profile evolution phenomenologically similar to GPM J1839-10 but with an interpulse, extending the geometric parameter space. Also of interest would be ASKAP J1755-2527, recently confirmed to be a 1.16-hour intermittent LPT [37], the emission-state-switching 54 minute LPT ASKAP J1935+2148 [38] and CHIME J0630++25 [39], whose 421 s period makes it an interesting counterpart to J1912-44.
Our results demonstrate that LPTs like GPM J1839-10 can be understood within the same emission geometry framework as radio emitting WD binary pulsars such as J1912-44 and AR Sco. The diversity in the observed periods, pulse morphologies, and polarisation states arise from variations in orbital synchronisation and viewing geometry.
The possible evolutionary connection between these systems remains debated, and we cannot exclude them being at different evolutionary stages but with common properties and geometries allowing them to emit bright polarised radio emission. With enough samples, we expect their evolutionary stages to emerge, and the methods developed in this work will enable insight into binary WD and magnetic field evolution in a way difficult to select from optical surveys alone.
Acknowledgements
This research is supported by an Australian Government Research Training Program (RTP) Scholarship . N.H.-W. is the recipient of an Australian Research Council Future Fellowship (project number FT190100231). N.R. is supported by the European Research Council (ERC) via the Consolidator Grant “MAGNESIA” (No. 817661) and the Proof of Concept “DeepSpacePulse” (No. 101189496), by the Catalan grant SGR2021-01269 (PI: Graber/Rea), the Spanish grant ID2023-153099NA-I00 (PI: Coti Zelati), and by the program Unidad de Excelencia Maria de Maeztu CEX2020-001058-M.
This scientific work uses data obtained from Inyarrimanha Ilgari Bundara, the CSIRO Murchison Radio-astronomy Observatory. We acknowledge the Wajarri Yamaji People as the Traditional Owners and native title holders of the Observatory site. CSIRO’s ASKAP radio telescope is part of the Australia Telescope National Facility (https://ror.org/05qajvd42). Operation of ASKAP is funded by the Australian Government with support from the National Collaborative Research Infrastructure Strategy. ASKAP uses the resources of the Pawsey Supercomputing Research Centre. Establishment of ASKAP, Inyarrimanha Ilgari Bundara, the CSIRO Murchison Radio-astronomy Observatory and the Pawsey Supercomputing Research Centre are initiatives of the Australian Government, with support from the Government of Western Australia and the Science and Industry Endowment Fund.
The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc.
We acknowledge Ben Stappers, Ewan Barr, and Yunpeng Men for contributing to the MeerKAT proposal. We acknowledge Scott Hyman, Tracy Clarke, and Simona Giacintucci for contributing to the VLA proposal. We thank Ingrid Pelisoli and Simone Scaringi for useful discussions on white dwarf evolution.
Author contributions
Cs.H performed the orbital timing, spin/beat period analysis, polarisation analysis, developed the geometric model, and contributed to the discussion in the manuscript. N.R contributed to the formulation of the model and its physical interpretation, and contributed to the discussion in the manuscript. N.H.-W led the proposals for the radio data toward GPM J1839-10, imaged and produced dynamic spectra from the radio data used in this work, and contributed to the formulation of the model and discussion in the manuscript. S.M contributed to the formulation of the model, and contributed to the discussion in the manuscript. R.P generated the observing scripts and calibrated the data for the VLA observations. E.L processed the ASKAP data and contributed to the polarisation analysis.
Competing interests
The authors declare no competing interests.
Data availability
Data supporting the findings of this study are available from the corresponding author upon request.
Code availability
The code used for the geometric model and MCMC fitting is available at
https://github.com/CsanadHorvath/GeometricLPT.git.
An interactive web visualisation is available at
https://www.desmos.com/3d/6bisspujwe.